A druggist mixed 100 cc of a 15% alcohol solution with 50 cc of pure alcohol. Find the percent concentration of the resulting mixture. Round to the nearest tenth of a percent.
Answer by checkley77(12844)

100*.15+1*50=150X
15+50=150X
150X=65
X=65/150
X=.43 OR 43% SOLUTION.